OpenStudy (anonymous):

the ratio 96000 miles to5600 in simplest form is

OpenStudy (ageta):

When I simplify, I get (96000 miles) / (5600 gallons) = (17.142857.. miles) / (1 gallon), or, in more common language, 17.142857.. miles

OpenStudy (anonymous):

7to12 or mabye i might be mixed and its 12to7

OpenStudy (anonymous):

http://www.hackmath.net/en/calculator/fraction?input=96000%2F5600&submit=Calculate 96000/5600 = 120⁄7 = 17+ 1⁄7 ≅ 17.14286